OpenStudy (anonymous):

What is a solution for the following system of equations? y=x^2 y=-4x+12 thank you.

OpenStudy (theviper):

Put the value of y in eq 2, i.e., \[\LARGE{y=-4x+12}\]\[\LARGE{x^2=-4x+12}\]\[\LARGE{\implies x^2+4x-12=0}\]

OpenStudy (theviper):

Now solve it by Quadratic Formula:)\[\Huge{\color{green}{-b \pm \sqrt{b^2-4ac}\over2a}}\]Can u do it now??

OpenStudy (anonymous):

Okay, Give me a moment to try this

OpenStudy (theviper):

a=1, b=4, c=-12.

OpenStudy (anonymous):

Okay. So \[-4\sqrt{4}^2-4(1)(-12) \]

OpenStudy (anonymous):

then divide by 4?

OpenStudy (anonymous):

The answer i got was -4

OpenStudy (anonymous):

i have to eat dinner. ill be right back!

OpenStudy (theviper):

Answer will be:- \[\LARGE{-4\pm \sqrt{4^2-4(1)(-12)}\over2(1)}\]

OpenStudy (theviper):

ok